Assume the following balance sheet for Bank X. Assume all numbers are in millions of
dollars. The required reserve ratio is 20%
Assets Liabilities
Reserves 102 Deposits 300
Securities 68
Loans 130
a) What is the amount of required reserves?
b) What is the maximum amount that the bank can lend?
c) If the central bank decided to sell 25 million dollar worth of government securities to bank X,
how will this transaction affect the bank’s balance sheet? What is the balance of securities and
required reserves and excess reserves after this transaction happens?
d) Assume the deposits increase by 10 million dollars. How will this affect the balance sheet?
Will the reserves, securities and loans change?
1
Expert's answer
2018-11-08T11:32:09-0500

rr = 20%

Assets Liabilities

Reserves 102 Deposits 300

Securities 68

Loans 130

a) The amount of required reserves RR = rr*D = 300*0.2 = 60.

b) The maximum amount that the bank can lend is 300 - 60 = 240.

c) If the central bank decided to sell 25 million dollar worth of government securities to bank X, then Securities = 68 + 25 = 93, RR = 60, ER = 102 - 60 - 25 = 17.

d) If the deposits increase by 10 million dollars, then the reserves will increase by 10*0.2 = 2, securities will remain the same and loans will increase by 10 - 2 = 8.